OpenStudy (anonymous):

What is the sum of the geometric sequence 2, 8, 32, … if there are 8 terms? @kc_kennylau

OpenStudy (kc_kennylau):

1. Find its common ratio

OpenStudy (kc_kennylau):

2. Apply S=a(r^n-1)/(r-1)

OpenStudy (anonymous):

thats 4 right?

OpenStudy (kc_kennylau):

yep :)

OpenStudy (kc_kennylau):

Then apply S=a(r^n-1)/(r-1) where a is the first term, r is the common ratio and n is the number of terms :D

OpenStudy (anonymous):

10922.66

OpenStudy (anonymous):

*cough* 43,690

OpenStudy (anonymous):

:( how @kc_kennylau

OpenStudy (kc_kennylau):

sorry back

OpenStudy (kc_kennylau):

mind showing me what you got?

OpenStudy (anonymous):

10922.66

OpenStudy (kc_kennylau):

I mean after you plugged it in, before calculating its value.

OpenStudy (kc_kennylau):

wait, I know what you did wrong

OpenStudy (anonymous):

ok \[2(4^{8-1})/4-1\]

OpenStudy (anonymous):

this is how i solved it

OpenStudy (kc_kennylau):

It's r^n-1 not r^(n-1) :)

OpenStudy (kc_kennylau):

PEMDAS P-parentheses E-exponents MD-multiplication+division AS-addition+subtraction

OpenStudy (anonymous):

699050.66

OpenStudy (anonymous):

so wait how does the equation look like

OpenStudy (kc_kennylau):

no, you must do it yourself, I cannot give you the answer

OpenStudy (kc_kennylau):

a(r^n-1)/(r-1) \[\Large\dfrac{a(r^n-1)}{r-1}\]

OpenStudy (kc_kennylau):

Just plug in those darn values

OpenStudy (anonymous):

sorry :(

OpenStudy (kc_kennylau):

not your fault :)

OpenStudy (anonymous):

43690? finger crossed

OpenStudy (kc_kennylau):

finally......

OpenStudy (kc_kennylau):

*drumroll*

OpenStudy (kc_kennylau):

BINGO :DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

lol ;) thanks

OpenStudy (kc_kennylau):

no problem :)